OpenStudy (anonymous):

What is the product of (-a + 3)(a + 4)? a^2 - a - 12 -a^2 - a - 12 -a^2 - a + 12 a^2 - a + 12

OpenStudy (anonymous):

C

OpenStudy (anonymous):

(x+9)(x+4)

OpenStudy (anonymous):

Factor a^2 - a - 20

OpenStudy (owlcoffee):

In order to find the solution of this, we'll have to take a look at the "distributive" postulate, wich fortunately for me, it's a postulate and it does not require proof. I'll show it to you: \[a(b+c)= ab+ac\] simple, it just states, that if we have a number multiplying the sum of other two, it's the same as multiplying each term by the number "a" and adding them: so for the problem: \[(-a+3)(a+4)= -a ^{2}-4a+3a+12\] simplifyin we get: \[-a ^{2}-a+12\]

OpenStudy (anonymous):

(a+4)(a-5) , @xo_kansasprincess_xo

OpenStudy (anonymous):

Factor x^2 - 5x - 14